### v1

# Rearrangements of distributions on integers that minimize variance

Abstract
Which permutations of a probability distribution on integers minimize variance?
Let $X$ be a random variable on a set of integers $\{x_{1},\dots,x_{N}\}$ such that $\mathbb{P}(X_{i}=x_{i})=p_{i}$, $i\in\{1,\dots,N\}$. Let $(p^{(1)},\dots,p^{(N)})$ be the sequence $(p_{1},\dots,p_{N})$ ordered non-increasingly. Let $X^{+}$ be the random variable defined by $\mathbb{P}(X=0)=p^{(1)}$, $\mathbb{P}(X=1)=p^{(2)}$, $\mathbb{P}(X=-1)=p^{(3)},\dots,\mathbb{P}(X=(-1)^{N}\lfloor\frac{N}{2}\rfloor)=p^{(N)}$. In this short note we generalize and prove the inequality ${\mathrm{Var}}\,X^{+}\leq{\mathrm{Var}}\,X$.
1 Introduction
Rearrangement inequalities, classically covered in Chapter X of Hardy, Littlewood and Pólya [1] have been applied to derive many other results, including isoperimetric inequalities, see, e.g., [7], and concentration function inequalities / variations of the Littlewood–Offord problem, see, e.g., [2, 3, 4, 5]. Many of the latter results have a form similar to the following one. Let $X_{1},\dots,X_{n}$ be independent random variables supported on finite sets of integers, and let $X_{1}^{+},\dots,X_{n}^{+}$ be independent random variables with the corresponding rearranged distribution functions. Then there exist $a_{1},\dots,a_{n}\in\{-1,1\}$ such that
$\max_{x\in\mathbb{Z}}\mathbb{P}(X_{1}+\dots+X_{n}=x)\leq\max_{x\in\mathbb{Z}}\mathbb{P}(a_{1}X_{1}^{+}+\dots+a_{n}X_{n}^{+}=x).$ (1)
For example, Theorem 371 of [1] implies that (1) holds (with $a_{1}=1$, $a_{2}=-1$ and $a_{3}=a_{4}=\dots=1$) in the case when $X_{i}^{+}$ are symmetric for all $i\geq 3$ and the main result of [4] is that (1) holds when $X_{i}$ is distributed uniformly on a finite subset of $\mathbb{Z}$ (in this case the signs $a_{i}$ are not important).
Consider another particular case where $X_{1},X_{2},\dots$ are i.i.d. copies of an integer random variable $X$ with a finite support, and assume that the support of XkX-k is not contained in $s\mathbb{Z}$ for some integers $k$ and $s$, $s>1$. In this case the local limit theorem, see, e.g., Theorem 1 in Chapter VII of [6], implies that
$\max_{x\in\mathbb{Z}}\mathbb{P}(X_{1}+\dots+X_{n}=x)=\frac{1+o(1)}{\sqrt{2\pi n{\mathrm{Var}}\,X}}$
and so (1) holds for $n$ large enough (with $a_{1}=\dots=a_{n}=1$) if
${\mathrm{Var}}\,X^{+}\leq{\mathrm{Var}}\,X$ (2)
and if the equality in (2) is only achieved in the obvious cases when $X-k\sim X^{+}$ or $X-k\sim-X^{+}$ for some integer $k$.
The question whether (2) always holds arose while applying a similar argument in [3]. In the present short note we provide a straightforward proof of (2) as we were not able to find it mentioned in the literature.
Let $f$ be the density function of an absolutely continuous random variable. $f$ can be transformed, see Chapter 10.12 of [1], to obtain a density $f^{*}$ called the symmetric decreasing rearrangement of $f$ which satisfies for any Borel set $B$ and the Lebesgue measure $\lambda$
$\int_{[-\frac{|B|}{2},\frac{|B|}{2}]}f^{*}d\lambda\geq\int_{B}fd\lambda.$ (3)
As for any non-negative random variable ${\mathbb{E}}\,X=\int_{t=0}^{\infty}\mathbb{P}(X>t)dt$, for any $p\geq 1$ we have ${\mathbb{E}}\,|X-{\mathbb{E}}\,X|^{p}=\int_{t=0}^{\infty}pt^{p-1}\mathbb{P}(|X-{\mathbb{E}}\,X|>t)dt$. If random variables $X$ and $X^{*}$ have densities $f$ and $f^{*}$ respectively, (3) implies that $\mathbb{P}(|X^{*}|>t)\leq\mathbb{P}(|X-{\mathbb{E}}\,X|>t)$ for any $t\geq 0$, so ${\mathbb{E}}\,|X^{*}|^{p}\leq{\mathbb{E}}\,|X-{\mathbb{E}}\,X|^{p}$. Thus, a ‘continuous’ variant of (2), as opposed to the integer variant that we consider here, follows rather easily, and has been noted in the literature, see, e.g. [8].
We will use the next definition.
Definition 1.1
Let $X$ be a random variable. Let $f:[0,+\infty)\to[0,+\infty)$ be a non-decreasing function such that $\inf_{a\in\mathbb{R}}{\mathbb{E}}\,f(|X-a|)<\infty$. Define a number
$D_{f}(X):=\min_{a\in\mathbb{R}}{\mathbb{E}}\,f(|X-a|).$
And the set
$M_{f}(X):=\arg\min_{a\in\mathbb{R}}{\mathbb{E}}\,f(|X-a|).$
Thus every $f$ as above gives a measure of dispersion $D_{f}$ and a central tendency $M_{f}$. These statistics can also be generalized to the $d$-dimensional Euclidean space or other normed spaces.
Theorem 1.2
Let $X$ be a random variable supported on a finite set of integers. Assume that $f:[0,+\infty]\to[0,+\infty)$ is non-decreasing. Then
$D_{f}(X^{+})\leq D_{f}(X).$ (4)
Furthermore, suppose that $f(x)$ has a positive derivative for $x>0$ and a right derivative at 0 such that $f^{\prime}(0+)=0$. Then (4) is strict unless XkX-k is distributed as $X^{+}$ or $X^{+}$ for some integer $k$.
Recall that $m$ is a median of $X$ if $\mathbb{P}(X\geq m)\geq\frac{1}{2}$ and $\mathbb{P}(X\leq m)\geq\frac{1}{2}$.
Corollary 1.3
Let $X$ and $f$ be as in Theorem 1.2.
  1. 1)
    If $f(x)=x$ for $x\geq 0$ then each value $m\in M_{f}(X)$ is a median of $X$ and $D_{f}(X)={\rm MAD_{median}}(X)={\mathbb{E}}\,|X-m|$, i.e., $D_{f}(X)$ is the mean absolute deviation of $X$ around the median.
  2. 2)
    If $f(x)=x^{2}$ for $x\geq 0$ then $M_{f}(X)=\{{\mathbb{E}}\,X\}$ and $D_{f}(X)={\mathrm{Var}}\,X$.
Thus ${\rm MAD_{median}}(X^{+})\leq{\rm MAD_{median}}(X)$ and ${\mathrm{Var}}\,X^{+}\leq{\mathrm{Var}}\,X$. Furthermore ${\mathrm{Var}}\,X^{+}={\mathrm{Var}}\,X$ if and only if $X-k\sim X^{+}$ or $X-k\sim-X^{+}$ for some integer $k$.
2 Proofs
Proof of Theorem 1.2 Let $x_{1},\dots,x_{N}$ be the support of $X$ listed in such a way that the corresponding probabilities $p^{(1)},\dots,p^{(N)}$ , given by $p^{(i)}=\mathbb{P}(X=x_{i})$, are non-increasing.
Let $a\in M_{f}(X)$. We denote $D_{f}(X)={\bf p}\cdot{\bf v}$ where ${\bf p}=(p^{(1)},p^{(2)},\ldots,p^{(N)})$ and
${\bf v}=(f(|x_{1}-a|),f(|x_{2}-a|),\ldots,f(|x_{N}-a|)).$
Let ${\bf{v^{\prime}}}=(f(|x^{(1)}-a|),f(|x^{(2)}-a|),\ldots,f(|x^{(N)}-a|))$ be the sequence $(f(|x_{1}-a|),f(|x_{2}-a|),\ldots,f(|x_{N}-a|))$ ordered non-decreasingly. Then, a classical result about the rearrangements of two sequences (e.g. Theorem 368 of [1]) implies that
${\bf p}\cdot{\bf v}\geq{\bf p}\cdot{\bf v^{\prime}}.$
Set $a^{\prime}=\min(a-\lfloor a\rfloor,\lfloor a\rfloor+1-a)$. In other words, the number $a^{\prime}\in[0,\frac{1}{2}]$ represents the distance between the number $a$ and its nearest integer. Set
${\bf w}=(f(a^{\prime}),f(1-a^{\prime}),f(1+a^{\prime}),f(2-a^{\prime}),f(2+a^{\prime}),\ldots,f(\lfloor\frac{N}{2}\rfloor+(-1)^{N-1}a^{\prime})).$
Clearly, $\bf w$} is ordered non-decreasingly. Further, recalling that $\{x^{(1)},\ldots,x^{(N)}\}$ is a set of NN distinct integers and $f$ is non-decreasing, it is not hard to see that every component of the vector $\bf{v^{\prime}}-\bf{w}$ is non-negative. Hence, we obtain that
${\bf p}\cdot{\bf v^{\prime}}\geq{\bf p}\cdot{\bf w}.$
Adding all the ingredients together we conclude that
$D_{f}(X)$ $={\bf p}\cdot{\bf v}$
$\geq{\bf p}\cdot{\bf v^{\prime}}$
$\geq{\bf p}\cdot{\bf w}$ (5)
$={\mathbb{E}}\,f(|X^{+}-a^{\prime}|)$ (6)
$\geq D_{f}(X^{+}).$
This finishes the proof of (4).
Assume now the additional properties of $f$ stated in the second part of the theorem. Assume $D_{f}(X)=D_{f}(X^{+})$, but $X$ is not a translation of $X^{+}$ or $X^{+}$. We will follow the proof of (4) and obtain a contradiction.
Since translating by a constant does not change $D_{f}(X)$, we can assume without loss of generality that $\lfloor a\rfloor=0$, equivalently, $a\in[0,1)$.
When defining $\bf v$ and $x_{1},\dots,x_{N}$ we may additionally assume that $(\mathbb{P}(X=x_{1}),-f(|x_{1}-a|))$, $\dots$, $(\mathbb{P}(X=x_{N}),-f(|x_{N}-a|))$ is ordered non-increasingly in lexicographic order.
We claim that
${\bf v}={\bf v^{\prime}}={\bf w}.$ (7)
To see the first equality, assume there exist $i$ and $j$ such that $i<j$ and $v_{i}>v_{j}$. Then due to to the ordering of $(x_{i})$, it must be $p^{(i)}>p^{(j)}$. This implies that $p^{(i)}v_{j}+p^{(j)}v_{i}<p^{(i)}v_{i}+p^{(j)}v_{j}$, so exchanging the atoms at $i$ and $j$ gives a random variable $X^{\prime}$, with $D_{f}(X^{\prime})<D_{f}(X)$, which is a contradiction to (4).
To see the second equality of (7), notice that since both of these vectors are ordered non-decreasingly, if they are not equal, we must have that some component of $\bf{v^{\prime}}-\bf{w}$ is positive, and hence (5) is strict, again a contradiction to (4).
Suppose first that $a^{\prime}\not\in\{0,\frac{1}{2}\}$. Then, since $f$ is strictly increasing for $x>0$, identity is the unique permutation that orders $(x_{i})$ non-decreasingly. When $a\in(0,\frac{1}{2})$ this corresponds to placing the probabilities $p^{(1)},\dots,p^{(N)}$ on $0,1,-1,\dots$ respectively as in the distribution of $X^{+}$. Similarly, when $a\in(\frac{1}{2},1)$, this corresponds to placing them on $1,0,2,-1,\dots$ respectively as in the distribution of $1-X^{+}$.
So we can assume that $a^{\prime}\in\{0,\frac{1}{2}\}$. Then, if $a^{\prime}=0$ we have $|x_{2k}|=|x_{2k+1}|$ for $k\in\{1,2,\dots\}$, and if $a^{\prime}=\frac{1}{2}$ we have $|x_{2k-1}-a^{\prime}|=|x_{2k}-a^{\prime}|$ for $k\in\{1,2,\dots\}$. It cannot be that for $a^{\prime}=0$ we have
$p^{(2k)}=p^{(2k+1)}\mbox{ for }k\in\{1,2,\dots\}$ (8)
or for $a^{\prime}=\frac{1}{2}$ we have
$p^{(2k-1)}=p^{(2k)}\mbox{ for }k\in\{1,2,\dots\}$ (9)
since in these cases (7) implies that $X\sim X^{+}$ (the distribution is symmetric around $a^{\prime}$).
Suppose that $a^{\prime}=a=0$. By the definition of $X^{+}$ we have $\mathbb{P}(X^{+}=k)\geq\mathbb{P}(X^{+}=-k)$ for all $k\in\{1,2,\dots\}$. Since (8) cannot hold, for some $k$ we have $\mathbb{P}(X^{+}=k)>\mathbb{P}(X^{+}=-k)$. Consider the function $g(x)={\mathbb{E}}\,f(|X^{+}-x|)$. By the assumptions on $f^{\prime}$ of the theorem, we have
$g^{\prime}(0+)=\mathbb{P}(X^{+}=0)f^{\prime}(0+)-\sum_{k\in\mathbb{Z}\setminus\{0\}}{\rm sgn}(k)\mathbb{P}(X^{+}=k)f^{\prime}(k)$
$=-\sum_{k\in\{1,2,\dots\}}(\mathbb{P}(X^{+}=k)-\mathbb{P}(X^{+}=-k))f^{\prime}(k)<0,$
so $D_{f}(X^{+})\leq g(\delta)<g(0)\leq D_{f}(X)$ for some $\delta>0$, a contradiction.
Finally, suppose that $a^{\prime}=a=\frac{1}{2}$. Note that by the definition of $X^{+}$, $\mathbb{P}(X^{+}=1-k)\geq\mathbb{P}(X^{+}=k)$ for $k\in\{1,2,\dots\}$. Since (9) cannot hold, for some $k$ we have $\mathbb{P}(X^{+}=1-k)>\mathbb{P}(X^{+}=k)$. Similarly as above
$g^{\prime}\left(\frac{1}{2}\right)=\sum_{k\in\{1,2,\dots\}}(\mathbb{P}(X^{+}=1-k)-\mathbb{P}(X^{+}=k))f^{\prime}\left(k-\frac{1}{2}\right)>0,$
so $D_{f}(X^{+})\leq g(\frac{1}{2}-\delta)<g(\frac{1}{2})={\mathbb{E}}\,f(|X^{+}-\frac{1}{2}|)\leq D_{f}(X)$ for some $\delta>0$, again a contradiction. $\Box$
Proof of Corollary 1.3 1) and 2) are folklore facts in statistics with straightforward proofs, see, e. g., Chapter 6 of [9]. The conclusion follows by applying Theorem 1.2. Note that in 2) we have $f^{\prime}(x)=2x>0$ for $x>0$ and $f^{\prime}(0+)=0$ as required. $\Box$

### v2

@@
 Let $(p^{(1)},\dots,p^{(N)})$ be the sequence $(p_{1},\dots,p_{N})$ ordered non-increasingly.
-Let $X^{+}$ be the random variable defined by $\mathbb{P}(X=0)=p^{(1)}$, $\mathbb{P}(X=1)=p^{(2)}$, $\mathbb{P}(X=-1)=p^{(3)},\dots,\mathbb{P}(X=(-1)^{N}\lfloor\frac{N}{2}\rfloor)=p^{(N)}$.
+Let $X^{+}$ be the random variable defined by $\mathbb{P}(X^{+}=0)=p^{(1)}$, $\mathbb{P}(X^{+}=1)=p^{(2)}$, $\mathbb{P}(X^{+}=-1)=p^{(3)},\dots,\mathbb{P}(X^{+}=(-1)^{N}\lfloor\frac{N}{2}\rfloor)=p^{(N)}$.
 In this short note we generalize and prove the inequality ${\mathrm{Var}}\,X^{+}\leq{\mathrm{Var}}\,X$.
@@
 In the present short note we provide a straightforward proof of (2) as we were not able to find it mentioned in the literature.
+A special case of the present result is used, along with many other ideas, in the proof of a much more general result of [3].
 Let $f$ be the density function of an absolutely continuous random variable.
@@
 Definition 1.1
-Let $X$ be a random variable. Let $f:[0,+\infty)\to[0,+\infty)$ be a non-decreasing function such that $\inf_{a\in\mathbb{R}}{\mathbb{E}}\,f(|X-a|)<\infty$. Define a number
-$D_{f}(X):=\min_{a\in\mathbb{R}}{\mathbb{E}}\,f(|X-a|).$
-And the set
-$M_{f}(X):=\arg\min_{a\in\mathbb{R}}{\mathbb{E}}\,f(|X-a|).$
+Let $X$ be a random variable. Let $f:[0,+\infty)\to[0,+\infty)$ be a non-decreasing function. Define a number
+$D_{f}(X):=\inf_{a\in\mathbb{R}}{\mathbb{E}}\,f(|X-a|)$
+and, in the case $D_{f}(X)<\infty$, the set
+$M_{f}(X):=\{a\in\mathbb{R}:{\mathbb{E}}\,f(|X-a|)=D_{f}(X)\}.$
 Thus every $f$ as above gives a measure of dispersion $D_{f}$ and a central tendency $M_{f}$. These statistics can also be generalized to the $d$-dimensional Euclidean space or other normed spaces.
 Theorem 1.2
-Let $X$ be a random variable supported on a finite set of integers. Assume that $f:[0,+\infty]\to[0,+\infty)$ is non-decreasing. Then
+Let $X$ be a random variable supported on a finite set of integers. Assume that $f:[0,+\infty)\to[0,+\infty)$ is non-decreasing and $D_{f}(X)<\infty$. Then
 $D_{f}(X^{+})\leq D_{f}(X).$ (4)
-Furthermore, suppose that $f(x)$ has a positive derivative for $x>0$ and a right derivative at 0 such that $f^{\prime}(0+)=0$.
+Furthermore, suppose that $f$ is continuous with a positive derivative for $x>0$ and a right derivative at 0 such that $f^{\prime}(0+)=0$.
 Then (4) is strict unless XkX-k is distributed as $X^{+}$ or $X^{+}$ for some integer $k$.
@@
 Proof of Theorem 1.2 Let $x_{1},\dots,x_{N}$ be the support of $X$ listed in such a way that the corresponding probabilities $p^{(1)},\dots,p^{(N)}$ , given by $p^{(i)}=\mathbb{P}(X=x_{i})$, are non-increasing.
-Let $a\in M_{f}(X)$. We denote $D_{f}(X)={\bf p}\cdot{\bf v}$ where ${\bf p}=(p^{(1)},p^{(2)},\ldots,p^{(N)})$ and
+It can be checked that $M_{f}(X)$ is non-empty when $f$ is continuous.
+First assume that $M_{f}(X)\neq\emptyset$.
+Let $a\in M_{f}(X)$.
+We denote $D_{f}(X)={\bf p}\cdot{\bf v}$ where ${\bf p}=(p^{(1)},p^{(2)},\ldots,p^{(N)})$ and
 ${\bf v}=(f(|x_{1}-a|),f(|x_{2}-a|),\ldots,f(|x_{N}-a|)).$
@@
 $\geq D_{f}(X^{+}).$
-This finishes the proof of (4).
+This finishes the proof of (4) when $M_{f}(X)\neq\emptyset$.
+For the general case, by definition, for any $\epsilon>0$ we can find $a=a(\epsilon)$ such that ${\mathbb{E}}\,f(|X-a|)\leq D_{f}(X)+\epsilon$.
+The same argument as above shows that $D_{f}(X)\geq D_{f}(X^{+})-\epsilon$.
+Since $\epsilon>0$ is arbitrary, we conclude that $D_{f}(X)\geq D_{f}(X^{+})$.
 Assume now the additional properties of $f$ stated in the second part of the theorem.
+Now $f$ is continuous, so $M_{f}(X)\neq\emptyset$.
 Assume $D_{f}(X)=D_{f}(X^{+})$, but $X$ is not a translation of $X^{+}$ or $X^{+}$.
@@
 Suppose first that $a^{\prime}\not\in\{0,\frac{1}{2}\}$.
-Then, since $f$ is strictly increasing for $x>0$, identity is the unique permutation that orders $(x_{i})$ non-decreasingly.
+Then, since $f$ is strictly increasing for $x>0$, identity is the unique permutation that orders the components of $\bf v$ non-decreasingly.
 When $a\in(0,\frac{1}{2})$ this corresponds to placing the probabilities $p^{(1)},\dots,p^{(N)}$ on $0,1,-1,\dots$ respectively as in the distribution of $X^{+}$.